OpenStudy (anonymous):

can someone help me with this equation? Multiply. State any excluded values. q+5/-3*-3q/q+2

OpenStudy (phi):

can you use the equation editor (see button below where you type in) or draw it?

OpenStudy (anonymous):

\[\frac{ q+5 }{ -3 }\times \frac{ -3q }{ q+2 }\]

OpenStudy (anonymous):

\[(-3q ^{2} - 15q) / (-3q - 6) = (-q ^{2} - 5q) / (-q - 2)\]

OpenStudy (phi):

You can't do much with it except "cancel" -3 from the top and bottom to get \[ \frac{q(q+5)}{q+2 } \] Of course, you do not want q= -2 because that would make the bottom 0 (-2+2=0) which is very bad. So you want q≠ -2

OpenStudy (anonymous):

thanks. i guess i'll give the other ones a try.